### v2

# Stable Gradients for Stable Learning at Scale in Deep Reinforcement Learning

Abstract
Scaling deep reinforcement learning networks is challenging and often results in degraded performance, yet the root causes of this failure mode remain poorly understood. Several recent works have proposed mechanisms to address this, but they are often complex and fail to highlight the causes underlying this difficulty. In this work, we conduct a series of empirical analyses which suggest that the combination of non-stationarity with gradient pathologies, due to suboptimal architectural choices, underlie the challenges of scale. We propose a series of direct interventions that stabilize gradient flow, enabling robust performance across a range of network depths and widths. Our interventions are simple to implement and compatible with well-established algorithms, and result in an effective mechanism that enables strong performance even at large scales. We validate our findings on a variety of agents and suites of environments. Source code here.
“We must be able to look at the world and see it as a dynamic process, not a static picture.”
— David Bohm
1 Introduction
Recent advances in deep reinforcement learning (deep RL) have demonstrated the ability of deep neural networks to solve complex decision-making tasks from robotics to game play and resource optimization (Mnih et al., 2015; Vinyals et al., 2019; Bellemare et al., 2020; Fawzi et al., 2022). Motivated by successes in supervised and generative learning, recent works have explored scaling architectures in deep RL, showing gains in representation quality and generalization across tasks (Farebrother et al., 2023; Taiga et al., 2023). However, scaling neural networks in deep RL remains fundamentally challenging (Ceron et al., 2024c, b). A central cause of this instability lies in the unique optimization challenges of RL. Unlike supervised learning, where data distributions are fixed, deep RL involves policy-dependent data that constantly change during training (Lyle et al., 2022). Each update of the policy $\pi_{\theta}$ alters future states and rewards, making the training objective inherently non-stationary. Value-based methods exacerbate these issues via bootstrapping, recursively using predicted values as targets.
Estimation errors compound over time (Fujimoto et al., 2018), especially under sparse or delayed rewards (Zheng et al., 2018), leading to unstable updates, policy collapse, or value divergence (Van Hasselt et al., 2016; Lyle et al., 2023, 2024). These dynamics are tightly coupled with architectural vulnerabilities. Deep networks face well known pathologies such as vanishing/exploding gradients (Pascanu et al., 2013), ill-conditioned Jacobians (Pennington et al., 2017), and activation saturation (Glorot and Bengio, 2010). In deep RL, these are magnified by the “deadly triad(Sutton and Barto, 2018; Van Hasselt et al., 2018), off-policy corrections, and changing targets. As networks scale, the risk of signal distortion and misalignment increases, resulting in underutilized capacity and brittle learning (Obando Ceron et al., 2023; Ceron et al., 2024b).
One overlooked source of these failures lies in how gradients propagate through the network. Specifically, the gradient decomposition, the layer-wise structure of backpropagation as a chain of Jacobians and weights determine how information flows during learning (Lee et al., 2020). While gradient signal preservation has been studied in supervised learning (Schoenholz et al., 2017; Jacot et al., 2018), its role in deep RL, where both inputs and targets shift continually, remains poorly understood.
In this work, we investigate how gradient decomposition interacts with non-stationarity and network scaling in deep RL. We demonstrate that in non-stationary settings like RL – where targets are bootstrapped, policies evolve continually, and data distributions shift – gradient signals progressively degrade across depth. This motivates the need for methods that explicitly preserve the structure of gradient information across layers. We explore this through a series of controlled experiments and ablations across multiple algorithms and environments, demonstrating that actively encouraging gradient propagation significantly improves stability and performance, even with large networks. Our work offers a promising approach for scaling deep RL architectures, yielding substantial performance gains across a variety of agents and training regimes.
2 Preliminaries
Deep Reinforcement Learning
A deep reinforcement learning agent interacts with an environment through sequences of actions ($a\in\mathcal{A}$), which produce corresponding sequences of observations ($s\in\mathcal{S}$) and rewards ($r\in\mathbb{R}$), resulting in trajectories of the form $\tau:=\{s_{0},a_{0},r_{0},s_{1},a_{1},r_{2},\ldots\}$. The agent’s behavior is often represented by a neural network with parameters $\theta$, composed of convolutional layers $\{\phi_{1},\phi_{2},\dots,\phi_{L_{c}}\}$ and dense (fully connected) layers $\{\psi_{1},\psi_{2},\dots,\psi_{L_{d}}\}$, where $\psi_{L_{d}}$ has an output dimensionality of $|\mathcal{A}|$. At every timestep $t$, an observation $s_{t}\in\mathcal{S}$ is fed through the network to obtain an estimate of the long-term value of each action: $Q_{\theta}(s_{t},\cdot)=\psi_{L_{d}}(\psi_{L_{d}-1}(\ldots(\phi_{L_{c}}(\ldots(\phi_{1}(s_{t}))\ldots))\ldots))$. The agent’s policy $\pi_{\theta}(\cdot\mid s_{t})$ specifies the probability of selecting each action, for instance by taking the softmax over the estimated values as in Eq. 1. The training objective is typically defined as the maximization of expected cumulative reward as in Eq. 2,
$\pi_{\theta}(a_{t}\mid s_{t})=\frac{e^{Q_{\theta}(s_{t},a_{t})}}{\sum_{a\in\mathcal{A}}e^{Q_{\theta}(s_{t},a)}}$ (1)
$J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$ (2)
where $\gamma\in[0,1)$ is a discount factor and $\tau$ denotes a trajectory generated by following policy $\pi_{\theta}$. Optimization proceeds by minimizing a surrogate loss $\mathcal{L}(\theta)$, which may be derived from temporal-difference (TD) errors, policy gradients, or actor-critic estimators (Sutton and Barto, 2018). In TD-based methods, the TD error at timestep $t$ is defined as:
$\delta_{t}=r_{t}+\gamma V_{\theta}(s_{t+1})-V_{\theta}(s_{t}),$
where $V_{\theta}(s)=\mathbb{E}_{a\sim\pi_{\theta}(a|s)}Q_{\theta}(s,a)$. The recurrent nature of $\delta_{t}$ introduces dependencies on both current estimates and future rewards, making $\mathcal{L}(\theta)$ inherently non-stationary. As the policy $\pi_{\theta}$ evolves, the data distribution used for training shifts, further complicating optimization. Training is performed by collecting trajectories, computing gradients $\nabla\mathcal{L}(\theta)$, and updating parameters via $\theta\leftarrow\theta-\eta\nabla\mathcal{L}(\theta)$, where $\eta>0$ is the learning rate. Following conventions from supervised learning, deep RL algorithms often use adaptive variants of stochastic gradient descent, such as Adam (Kingma and Ba, 2014) or RMSprop (Hinton, 2012), which adjust learning rates based on running estimates of gradient statistics. The gradients with respect to each layer are denoted by;
$\nabla\phi_{i}=\frac{\partial\mathcal{L}}{\partial\phi_{i}},\quad\nabla\psi_{j}=\frac{\partial\mathcal{L}}{\partial\psi_{j}},$
where $\phi_{i}$ and $\psi_{j}$ represent the parameters (i.e., weight matrices or bias vectors) of layer $i$ and $j$ respectively. The structure and magnitude of these gradients ($\nabla\phi_{i}$ and $\nabla\psi_{j}$) are influenced by the loss function, data distribution collected from the environment, and the architecture itself. These per-layer gradients determine how effectively different parts of the network adapt during training.
While training large models in supervised learning settings present challenges, advances in initialization, normalization, and scaling strategies have enabled relatively stable optimization (Ioffe and Szegedy, 2015; Ba et al., 2016; Glorot and Bengio, 2010). Scaling up model size has been a central driver of progress across domains, improving generalization, enhancing representation learning, and boosting downstream performance (Kaplan et al., 2020).
Deep RL differs substantially from supervised learning. First, the data distribution is non-stationary, continually shifting as $\pi_{\theta}$ updates. Second, learning signals are often sparse, delayed, or noisy, which introduces variance in the estimated gradients (Han et al., 2022; Fujimoto et al., 2018; Liu et al., 2025b). These factors destabilize optimization and lead to loss surfaces with sharp curvature and complex local structure (Ilyas et al., 2020; Achiam et al., 2019). Moreover, increasing model capacity often degrades performance unless regularization or architectural interventions are applied (Gogianu et al., 2021; Bjorck et al., 2021; Schwarzer et al., 2023; Wang et al., 2025).
These challenges are further compounded by both architectural and environmental factors. Network depth, width, initialization, and nonlinearity affect how gradients are propagated across layers. Meanwhile, reward sparsity, exploration dificulty, and transition stochasticity impose additional structure on the optimization landscape. The resulting geometry reflects the joint dynamics of policy, environment, and architecture, making deep RL optimization uniquely complex.
Gradient Propagation
Training deep networks poses fundamental challenges for effective gradient propagation (Glorot and Bengio, 2010). As network depth increases, gradients may either vanish or explode as they are backpropagated through multiple layers, impeding the optimization of early layers and destabilizing learning dynamics (Ba et al., 2016). These issues arise from repeated applications of the chain rule. For a network with intermediate hidden representations $\{h_{0},h_{1},\dots,h_{L}\}$, where $h_{k}\in\mathbb{R}^{d_{k}}$, the gradient of the loss $\mathcal{L}$ with respect to a hidden layer $h_{\ell}$ is:
$\frac{\partial\mathcal{L}}{\partial h_{\ell}}=\left(\prod_{k=\ell+1}^{L}\frac{\partial h_{k}}{\partial h_{k-1}}\right)\frac{\partial\mathcal{L}}{\partial h_{L}},$
where each $\frac{\partial h_{k}}{\partial h_{k-1}}\in\mathbb{R}^{d_{k}\times d_{k-1}}$ is the Jacobian. If the singular values of these Jacobians are not properly controlled, their repeated multiplication can cause the norm of the gradient to shrink or grow exponentially with $\mathcal{L}$. This severely impairs convergence, as earlier layers receive little to no useful gradient signal or become numerically unstable (Ioffe and Szegedy, 2015; He et al., 2016).
In addition to depth, the width of the network also influences gradient propagation. Consider a fully connected layer with weight matrix $W\in\mathbb{R}^{m\times n}$ and input vector $h\in\mathbb{R}^{n}$. The output is $Wh\in\mathbb{R}^{m}$, and under the assumption that $W$ and $h$ have i.i.d. zero-mean entries with finite variance $\sigma_{W}$ and $\sigma_{h}$, respectively, the variance of the output is given by $\mathrm{Var}[Wh]=n\,\sigma_{W}\,\sigma_{h}$. Thus, scaling the width $n$ without adjusting $\sigma_{W}$ and $\sigma_{h}$ leads to instability in forward and backward signal propagation affecting gradient norms and optimization trajectories.
Beyond depth and width, the choice of nonlinearity also plays a central role in determining how gradients propagate . In a typical feedforward network, hidden activations evolve as $h_{k}=\zeta(W_{k}h_{k-1})$, where $\zeta(\cdot)$ is a nonlinear activation function (e.g., ReLU, $\tanh$, sigmoid), and $W_{k}$ is the weight matrix at layer $k$. During backpropagation, the gradient with respect to a hidden layer includes the product of the Jacobian of the linear transformation and the derivative of the nonlinearity:
$\frac{\partial\mathcal{L}}{\partial h_{k-1}}=W_{k}^{\top}\left(\zeta^{\prime}(W_{k}h_{k-1})\odot\frac{\partial\mathcal{L}}{\partial h_{k}}\right),$
where $\zeta^{\prime}(\cdot)$ denotes the elementwise derivative of the activation function, and $\odot$ represents elementwise multiplication. For ReLU, $\zeta^{\prime}(x)=\mathbf{1}_{x>0}$, so the gradient is entirely blocked wherever the neuron is inactive. This leads to the well-known dying ReLU problem, where a significant portion of the network ceases to update and becomes untrainable (Lu et al., 2019; Shin and Karniadakis, 2020).
3 Diagnosis: Gradients Under Non-stationarity and Scale
A fundamental premise of modern deep learning is that scaling model capacity yields consistent gains in performance (Kaplan et al., 2020; Chowdhery et al., 2023). This has held true in large-scale supervised learning, where training data distributions are stationary and i.i.d., and gradient descent operates under relatively stable conditions. However, in non-stationary settings, such as RL, gradient-based optimization faces severe challenges that scaling alone may exacerbate (Ceron et al., 2024b, c). In this section, we diagnose how gradient pathologies emerge and intensify across different settings, with a focus on architectural scaling in width and depth (network scales used specified in Table 1).
3.1 Gradient Pathologies
We train neural networks of varying depths and widths and analyze their training dynamics.
Supervised Learning (Stationary and Non-Stationary) We use the CIFAR-10 image classification benchmark (Krizhevsky and others, 2009), where the input-output mapping remains fixed over time. Models consist of standard 6-layer convolutional neural networks (CNN) followed by a multi-layer perceptron (MLP). We vary the depth and width of the MLP to explore how model scale influences learning behavior. To introduce non-stationarity, we periodically shuffle the training labels during training, following the setup by Sokar et al. (2023). This creates a loss landscape that changes over time, echoing the challenges of deep RL. Fig. 1 illustrates the contrast in training behavior and gradient flow between stationary and non-stationary supervised learning. Under non-stationarity, deep networks fail to recover accuracy, which aligns with a marked degradation in gradient magnitudes.
height 122pt depth 0 pt width 1.3 pt Refer to caption
Figure 1: Training dynamics under stationary and non-stationary supervised learning. (Left) In the stationary setting, both shallow and deep models fit the data effectively across widths. Under non-stationarity only shallow networks partially recover during training, while deeper ones collapse. (Right) This collapse correlates with degraded gradient flow. In stationary settings, gradient norms remains stable across all network scales (shaded boxes) while in non-stationary settings (solid-colored boxes), gradient magnitudes diminish with depth and width, suggesting poor adaptability.
Reinforcement Learning As discussed in Sec. 2, RL introduces fundamentally different sources of non-stationarity due to the policy-dependent data distribution and moving target estimates. To study gradient dynamics, we use PQN (Gallici et al., 2025), a recent value-based algorithm that achieves strong performance without relying on a target network or replay buffer. PQN ensures stability and convergence using Layer Normalization (Ba et al., 2016) and supports GPU-based training through vectorized environments for online parallel data collection. In Sec. C.1 we extend our investigation to DQN (Mnih et al., 2015) and Rainbow (Hessel et al., 2018), demonstrating the generality of our observations. As shown in Fig. 2, deeper networks trained with PQN exhibit a collapse in both episode returns and gradient norms [^1] Unless otherwise specified, all ALE results are averaged over three seeds., highlighting the fragility of deep models under non-stationarity.
height 122pt depth 0 pt width 1.3 pt Refer to caption
Figure 2: Mean episode returns and gradient norms across increasing MLP depths and widths on two ALE games using PQN. (Left) Only shallow networks achieve high episode returns; performance collapses for deeper networks. (Right) The collapse correlates with vanishing gradient norms, suggesting that deeper models fail to adapt to non-stationarity in deep RL.
3.2 Training Degradation
In Fig. 3 we evaluate diagnostic metrics capturing expressivity and training dynamics, revealing that deeper networks exhibit pronounced training pathologies and degraded performance. We first measure the fraction of dormant neurons, defined as units with near-zero activations over a batch of trajectories (Sokar et al., 2023), and find that dormant neurons grow with depth, signaling underutilized capacity. Next, we assess representational expressivity using SRank, the effective rank of penultimate-layer activations (Kumar et al., 2020), observing that deeper networks tend to collapse state representations into lower-dimensional, and less expressive (as evidenced by declining returns) subspaces. To study loss curvature, we compute the Hessian trace of the temporal-difference loss. This metric serves as a proxy for sharpness or smoothness in optimization (Ghorbani et al., 2019), similarly to tracking the largest eigenvalue. Fig. 3 shows that only shallow networks exhibit high Hessian trace values, suggesting access to sharper regions of the loss surface with pronounced directions of improvement. In contrast, deeper architectures consistently show near-zero trace, indicating poorly conditioned geometry that hinders effective gradient-based updates. These findings suggest a breakdown in representation, plasticity, and optimization as networks scale, ultimately impeding learning.
Figure 3: Training pathologies emerge as MLP depth increases. Deeper networks exhibit a higher fraction of inactive neurons, reduced representation rank (SRank), vanishing Hessian trace (loss curvature), and degraded learning performance (mean Q-values and episode returns). These trends indicate that scaling depth limits expressivity and plasticity, impairing policy quality.
Key observations on gradients under non-stationarity and scale: Non-stationarity amplifies gradient degradation in deeper and wider networks. In deep RL, deeper models suffer from vanishing gradients, reduced activations, and loss of representational expressivity. The flat loss curvature intensifies with depth, correlating with poor learning.
4 Stabilizing Gradients
Having identified the pathologies that emerge in non-stationary regimes, particularly under large-scale architectures, we investigate strategies to mitigate these instabilities. We focus on two complementary interventions: skip connections (He et al., 2016) and optimizers (Martens and Grosse, 2015), as these directly improve gradient flow. We continue to use PQN as our base RL algorithm and evaluate on the Atari-10 suite (Aitchison et al., 2023). In Sec. 5, we demonstrate that the effectiveness of our proposed gradient interventions generalize beyond this specific algorithm and environment suite.
4.1 Intervention 1: Multi-Skip Residuals for Gradient Stability
Gradient instability in deep networks is often aggravated by increasing depth, non-linear activations, and misaligned curvature across layers. While standard residual connections offer some relief by introducing shortcut paths for gradient flow (He et al., 2016), they typically span only one or two layers, which can be insufficient in the presence of severe gradient disruption due to non-stationarity. We introduce multi-skip residual connections, in which the flattened convolutional features are broadcast directly to all subsequent MLP layers. This design ensures that gradients can propagate from any depth back to the shared encoder without obstruction.
We compare our network architecture against the standard fully connected baseline across varying depths. As shown in Fig. 5 (left), performance collapses with increased depth in the baseline, while the multi-skip architecture maintains stable learning and continues to improve across widths. This improvement is accompanied by consistently higher gradient magnitudes. Complete results across all network depths and widths are presented in Sec. C.4.
Figure 4: (Left) MLP architectures and (right) scaling strategies studied.
height 110pt depth 0 pt width 1.3 pt Refer to captionRefer to caption
Figure 5: Gradient-stabilizing interventions improve scalability in deep RL. (Left) Standard fully connected networks trained with PQN collapse at greater depths due to vanishing gradients. In contrast, multi-skip architectures maintain gradient flow and scale effectively. (Right) The default RAdam optimizer leads to instability in deep networks, while switching to the Kron optimizer preserves gradient signal and enables stable learning without architectural changes.
4.2 Intervention 2: Second-Order Optimizers for Non-Stationarity
First-order optimizers such as SGD and Adam rely on local gradient estimates and fixed heuristics (e.g., momentum, adaptive step sizes) (Kingma and Ba, 2014), which are agnostic to curvature and often brittle under shifting data distributions. In contrast, second-order methods adjust parameter updates using curvature information, enabling more informed and stable adaptation.
Let $\mathcal{L}(\theta)$ denote the loss function, and $g=\nabla\mathcal{L}(\theta)$ its gradient. A second-order update takes the form $\theta_{t+1}=\theta_{t}-\eta H^{-1}g$, where $H$ is the curvature matrix, typically the Hessian or the Fisher Information Matrix (FIM) (Martens, 2020). Directly inverting $H$ is computationally infeasible in deep neural networks so Kronecker-factored approximations, such as K-FAC (Martens and Grosse, 2015), address this challenge by approximating $H$ using low-rank Kronecker products.
Kronecker-factored optimizer (Kron for short) approximates the FIM and applies structured preconditioning that captures inter-parameter dependencies, unlike Adam’s diagonal scaling. This yields directionally aware preconditioning that better aligns with the curvature of the loss surface (Martens, 2020). In non-stationary settings, such as deep RL, where both the data distribution and curvature evolve over time, curvature-aware updates can help preserve gradient signal by maintaining stable update magnitudes and directions. As shown in Fig. 5 (right), replacing RAdam with Kron prevents performance collapse at greater depths, even in standard MLP architectures. Complete results across all network depths and widths are presented in Sec. C.6.
4.3 Combining Gradient Interventions
We combine both gradient interventions to PQN and evaluate it on the full ALE suite (57 games), across 3 seeds and 200M frames. Fig. 6 shows that our augmented agent outperforms the baseline in 90% of the environments, achieving a median relative improvement of 83.27%. Notably, the baseline PQN is itself competitive with strong agents such as Rainbow (Gallici et al., 2025), highlighting the effectiveness of our interventions. Detailed per-environment learning curves can be found in Sec. E.1.
In Fig. 7 we validate the effectiveness of the combined gradient interventions in the non-stationary SL setting we used as motivation in Sec. 3. The results verify that these interventions enable high accuracy and sustained adaptability across depths and widths, even under dynamic label reshuffling.
height 90pt depth 0 pt width 1.3 pt Refer to caption
Figure 6: Gradient-stabilized PQN achieves superior scalability. (Left) On Atari-10, the combined interventions lead to high HNS even at greater depths, outperforming either intervention alone (see Fig. 5) and increased gradient gradient flow. (Right) On the full ALE suite, our agent outperforms the baseline in 90% of the games with a median performance improvement of 83.27%.
height 122pt depth 0 pt width 1.3 pt Refer to caption
Figure 7: Gradient interventions enable rapid recovery in non-stationary SL. (Left) Models with combined gradient interventions rapidly recover accuracy after label reshuffling, demonstrating robust adaptation in non-stationary settings. (Right) This is supported by stable gradient flow across depth. Dashed curves and shaded boxes indicate MLP baselines.
4.4 Alternative Gradient-Stabilization Methods
To ensure that our findings were not specific to a narrow choice of interventions, we conducted a broader exploration of alternative strategies for improving gradient stability in deep RL. We tested a variety of approaches inspired by prior work on optimization and representation stability in both supervised and RL settings (see Sec. C.7 for more details).
As summarized in Tab. 3, none of these methods consistently improved stability or performance compared to our proposed combination of multi-skip residuals and Kronecker-factored optimization. In many cases, the alternatives yielded either negligible gains or degraded performance as network depth increased, reinforcing that architectural and curvature-aware interventions are key to preserving gradient flow at scale.
5 Beyond the ALE and PQN
To evaluate the generality of our findings, we extend our analyses. Specifically, we: (i) apply our proposed methods to PPO (Schulman et al., 2017) on the full ALE and on continuous control tasks in Isaac Gym (Makoviychuk et al., 2021); (ii) assess the impact of richer convolutional encoders by replacing the standard CNN backbone used in the ALE with the Impala CNN architecture (Espeholt et al., 2018); (iii) augment Simba (Lee et al., 2025) with our proposed techniques and evaluate performance on the DeepMind Control Suite (DMC) (Tassa et al., 2018); and (iv) investigate whether our interventions can stabilize and scale a wide range of Q-learning algorithms in challenging offline RL benchmarks (Park et al., 2025).
height 110pt depth 0 pt width 1.3 pt Refer to caption
Figure 8: PPO with gradient interventions. Left: On the full ALE suite, applying the combined gradient interventions to PPO yields a median performance improvement of 31.40% and outperforms the baseline in 83.64% of the games. Right: In the Cartpole and Anymal tasks from IsaacGym, only the augmented PPO maintains stable performance across depths and widths.
PPO with Gradient Interventions.
Fig. 8 (left) shows that augmenting PPO with the same strategies as in PQN (Layer Normalization by default on PQN, multi-skip residual connections, and Kronecker-factored optimization) significantly boots performance. On the ALE benchmark, the augmented PPO outperforms the baseline in 83.64% of the environments, achieving a median relative improvement of 31.40%. In Isaac Gym’s continuous control tasks, including Cartpole and Anymal (Fig. 8, right), the baseline PPO collapses as model size increases, while the augmented variant remains stable and achieves superior performance at all depths and widths.
Gradient Interventions in Scaled Encoder Variants
The Impala CNN is a scalable convolutional architecture that has demonstrated strong performance gains in agents such as Impala (Espeholt et al., 2018) and Rainbow (Hessel et al., 2018). We investigate whether, given its capacity to extract richer representations from visual input, combining Impala CNN with our gradient flow interventions enables effective scaling of the MLP component. As shown in Fig. 9, PPO and PQN benefit significantly from replacing the standard CNN with the Impala CNN. For PQN, the Impala encoder enables successful scaling of the MLP, in contrast to the performance collapse seen without our interventions. These results suggest that the expressivity of richer visual encoders is more effectively leveraged by deeper networks when gradient flow is preserved.
height 85pt depth 0 pt width 1.3 pt Refer to caption
Figure 9: Scaling performance with standard vs. Impala CNN encoders on PQN (left) and PPO (right). Each agent is evaluated using both the Atari CNN (left sub-panels) and the Impala CNN (right sub-panels) as the encoder. Gradient interventions enable successful scaling in both cases.
Simba with Kron Optimizer.
Simba (Lee et al., 2025) is a scalable actor-critic framework that integrates normalization, residual connections, and LayerNorm. We augment Simba by replacing its default AdamW optimizer with Kron while keeping all other hyperparameters fixed. We evaluate SAC (Haarnoja et al., 2018) and DDPG (Lillicrap et al., 2015) on challenging DMC tasks, using the Simba architectures of varying depth and width. Despite its design for scalability, default Simba collapses across all tasks as networks grow as shown in Fig. 10 (additional results in Sec. E.2). In contrast, the Kron-augmented version successfully scales in both depth and width, achieving consistent and stable performance gains. These findings underscore the generality of our approach as effectively enabling parameter scaling in deep RL agents.
Figure 10: Performance comparison between AdamW (dashed lines) and Kron (solid lines) optimizers using the SimBa architecture with SAC and DDPG, averaged over 55 random seeds. As model size increases, AdamW leads to consistent performance degradation, while Kron enables stable and improved learning with larger networks.
Gradient Interventions for Scalable Offline Q-Learning.
Park et al. (2025) highlight significant challenges in scaling Q-learning algorithms for complex offline RL tasks, demonstrating that many standard offline RL baselines fail to learn effective policies, even on large, high-quality datasets. Their key finding was that performance improvements were primarily driven by techniques that shorten the effective credit assignment horizon, such as n-step returns and hierarchical methods. This led to their proposal of two new high-performing algorithms, SHARSA and DSHARSA, which are designed to operate with shorter effective horizons. This finding motivates a parallel investigation: can our proposed gradient interventions, which are designed to stabilize and accelerate deep network training, also address the scaling limitations of offline Q-learning? To test this, we augment the full suite of baselines and novel algorithms from Park et al. (2025) with our proposed gradient interventions. The results, presented in Fig. 11, show that our methods provide a complementary path to scalability. Applying our interventions generally improves the performance of the baselines across all tasks.
The performance gains are particularly pronounced in the most sparse-reward task, humanoidmaze-giant-navigate, where our gradient interventions enable multiple methods to achieve near-optimal performance, whereas their baseline counterparts largely fail. Furthermore, this stability extends to generalization across task difficulty. When moving from puzzle-4x5-play to the harder puzzle-4x6-play task, many baselines exhibit a sharp performance degradation. In contrast, the performance of several algorithms with our interventions remains consistent and high, demonstrating improved robustness. Finally, we note that while the primary focus of this paper is to address gradient pathologies arising from scaling and non-stationarity, these results highlight that our interventions are also highly beneficial in offline deep RL, where inputs are stationary.
Figure 11: Performance of offline Q-learning algorithms with and without our gradient interventions. We compare the original algorithms from Park et al. (2025) against our augmented versions. The results, averaged over 33 seeds demonstrate a general improvement in scalability.
6 Related Work
A central challenge in scaling deep RL lies in the inefficient use of model capacity. Increasing parameter counts often fails to yield proportional gains due to under-utilization. Sokar et al. (2023); Mayor et al. (2025) and Liu et al. (2025c) show that online RL induces a growing fraction of inactive neurons, a phenomenon also observed in offline settings. Ceron et al. (2024b) report that up to 95% of parameters can be pruned post-training with negligible performance drop, underscoring substantial redundancy. These findings have motivated techniques such as weight resetting (Schwarzer et al., 2023), tokenized computation (Sokar et al., 2025), and sparse architectures (Ceron et al., 2024c; Willi et al., 2024; Liu et al., 2025a; Ma et al., 2025), along with auxiliary objectives to promote capacity utilization (Farebrother et al., 2023). While scaling model size offers greater expressivity, its benefits depend on appropriate training strategies (Ota et al., 2021). Architectural interventions such as SimBa (Lee et al., 2025) improve robustness by regularizing signal propagation through components such as observation normalization, residual feedforward blocks, and layer normalization. Complementarily, BRO (Nauman et al., 2024) shows that scaling the critic network yields substantial gains in sample and compute efficiency, provided it is paired with strong regularization and optimistic exploration strategies.
Gradient flow, however, remains a central bottleneck. We complement prior efforts by explicitly targeting vanishing gradients as a mechanism for improving scalability. Our approach builds on the role of LayerNorm in stabilizing training and enhancing plasticity (Lyle et al., 2024), and leverages its theoretical effect on gradient preservation as formalized in PQN (Gallici et al., 2025). Optimization-level interventions such as second-order methods (Martens and Grosse, 2015; Muppidi et al., 2024) and adaptive optimizers (Ellis et al., 2024; Bengio et al., 2021; Wu et al., 2017) also address instability under non-stationarity. Our approach integrates architectural and optimizer-level interventions to enable stable gradient flow and unlock parameter scaling in deep RL agents.
7 Discussion
Our analyses in Sec. 3 suggest that the difficulty in scaling networks in deep RL stems from the interaction between inherent non-stationarity and gradient pathologies that worsen with network size. In Sec. 4, we introduced targeted interventions to address these challenges, and in Sec. 4.3, we demonstrated their effectiveness. We validated the generality of our approach across agents and environment suites, consistently observing similar trends. These findings reaffirm the critical role of network design and optimization dynamics in training scalable RL agents. While our proposed solutions may not be optimal, they establish a strong baseline and provide a foundation for future work on gradient stabilization in deep RL. More broadly, our findings suggest that scaling limitations in deep RL are not solely attributable to algorithmic instability or insufficient exploration, but also stem from gradient pathologies amplified by architectural and optimization choices. Addressing these issues directly, without altering the learning algorithm, yields substantial gains in scalability and performance. This suggests that ensuring stable gradient flow is a necessary precondition for effective parameter scaling in deep RL.
Limitations. Our study is constrained by computational resources, which limited our ability to explore architectures beyond a certain size. While our interventions show consistent improvements across agents and environments, further scaling remains an open question. While using second order optimizers introduced additional computational overhead (see Tab. 12), this cost is mitigated by leveraging vectorized environments and efficient deep RL algorithms, narrowing the gap relative to standard methods. These limitations highlight promising directions for future work, including the development of more computationally efficient gradient stabilization strategies and scalable optimization techniques.
Acknowledgment
The authors would like to thank João Guilherme Madeira Araújo, Evan Walters, Olya Mastikhina, Dhruv Sreenivas, Ali Saheb Pasand, Ayoub Echchahed and Gandharv Patil for valuable discussions during the preparation of this work. João Araújo deserves a special mention for providing us valuable feed-back on an early draft of the paper. We want to acknowledge funding support from Natural Sciences and Engineering Research Council (NSERC) of Canada, Google Research, Fonds de recherche du Québec (FRQNT) and The Canadian Institute for Advanced Research (CIFAR) and compute support from Digital Research Alliance of Canada, Mila IDT, and NVidia. We would also like to thank the Python community Van Rossum and Drake Jr (1995); Oliphant (2007) for developing tools that enabled this work, including NumPy Harris et al. (2020), Matplotlib Hunter (2007), Jupyter Kluyver et al. (2016), and Pandas McKinney (2013).
Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Environment Details
Throughout the paper, we evaluate the deep reinforcement learning agents’ performance on the Atari-10 suite [Aitchison et al., 2023], a curated subset of games from the Arcade Learning Environment (ALE) [Bellemare et al., 2013]. Atari-10 consists of 10 games selected to capture the maximum variance in algorithm performance, achieving over 90% correlation with results on the full ALE benchmark. This makes it a computationally efficient yet representative testbed for deep reinforcement learning. We follow the experimental protocol of Obando Ceron et al. [2023], Ceron et al. [2024c], Agarwal et al. [2021], running each experiment with three random seeds and reporting the aggregate human-normalized score across games.
The games in Atari-10 are:
  • Amidar, Battle Zone, Bowling, Double Dunk, Frostbite, Kung Fu Master, Name This Game, Phoenix, Q*Bert and River Raid.
Additionally, to further support the generality of our findings, we evaluate the proposed combined gradient interventions on the full ALE benchmark. We also assess their effectiveness on continuous control tasks from the IsaacGym simulator [Makoviychuk et al., 2021] and the DeepMind Control Suite (DMC) [Tassa et al., 2018], extending our analysis to robotics-based environments. We conduct experiments on the 4 challenging tasks of DMC:
  • Humanoid Walk, Humanoid Run, Dog Trot and Dog Run.
Appendix B Network Sizes
Throughout the paper, we experiment with models of varying depths and widths. Unless stated otherwise (e.g. in  Sec. 5, where we evaluate the Impala CNN), the convolutional feature extractors are kept fixed. Consequently, our experiments focus primarily on scaling strategies and architectural variations in the MLP components of the networks.
To enable meaningful comparisons across different learning regimes, the MLP architectures are kept consistent across supervised learning (SL), non-stationary SL, and reinforcement learning (RL) experiments. This consistency ensures that observed differences in gradient behavior arise from the learning setting itself, rather than confounding factors due to domain-specific architectures.
Table 1 provides detailed information on the number of parameters for each depth–width configuration, categorized as small, medium, or large, as used throughout the paper.
Table 1: Number of parameters (in millions) for different MLP architectures.
[…]
Appendix C Additional Experiments
C.1 Scaling with DQN and Rainbow
To further support our hypothesis on the emergence of gradient pathologies in deep reinforcement learning, we investigate whether similar issues arise in algorithms beyond PQN and PPO, as discussed in the main paper. Specifically, we study the effects of architectural scaling on two widely used value-based algorithms: DQN [Mnih et al., 2015] and Rainbow [Hessel et al., 2018].
DQN is a foundational deep RL algorithm that learns action-value functions using temporal difference updates and experience replay, serving as a standard baseline for value-based methods. Rainbow extends DQN by integrating several enhancements, such as double Q-learning, prioritized experience replay, dueling networks, multi-step learning, distributional value functions, and noisy exploration, to achieve improved sample efficiency and stability.
In Fig. 12, we report the performance of DQN and Rainbow as we scale the depth and width of their networks. As with PQN and PPO, we observe consistent degradation in performance at larger scales. In Fig. 13, we present the corresponding gradient behavior, which reveals the same vanishing and destabilization phenomena discussed in this work. These findings reinforce the generality of the identified gradient pathologies across both policy-based and value-based deep RL algorithms.
(a) DQN HNS
(b) Rainbow HNS
Figure 12: Median human normalized scores for DQN (left) and Rainbow (right) as a function of total network parameters. Lines of different colors denote varying network depths, while marker shapes indicate different widths. For both agents, performance consistently declines as network size increases, highlighting the adverse effects of scaling.
Figure 13: Gradient magnitudes during training for DQN (top) and Rainbow (bottom). As network depth increases, gradient flow systematically diminishes, ultimately collapsing to near-zero values. This consistent decay mirrors the performance degradation observed at larger scales.
C.2 Combining Gradient Interventions in Non-stationary Supervised Learning
Building on our findings in Sec. 4.3, we extend our analysis by applying the proposed combined gradient interventions to the same image classification models used in Sec. 3. Specifically, we train the models in the non-stationary supervised learning setup, where the CIFAR-10 labels are iteratively shuffled, following the experimental design from Sokar et al. [2023]. As demonstrated in Sec. 3, while models in standard supervised learning settings are able to scale effectively and maintain high performance, introducing non-stationarity leads to failure in adaptation for baselines that use fully connected layers and the Adam optimizer. This issue is exacerbated as model scale increases.
Our results, presented in Fig. 7, show that combining the multi-skip architecture for the MLP component with the Kronecker-factored optimizer and Layer Normalization enables near-perfect continuous adaptation. The models quickly adapt to the changing optimization problem following label reshuffling, with gradient magnitudes remaining stable throughout the process.
C.3 Architecture and Optimizer Ablations
In this work, we introduce the multi-skip architecture, an extension of the standard residual MLP design, and propose the use of the Kronecker-factored optimizer for online deep RL. While these techniques form the basis of our primary interventions, our broader goal is not to prescribe a fixed set of methods, but rather to motivate a general class of architectural and optimization interventions that promote healthy gradient flow in deep networks. To this end, we expand the scope of our evaluation by incorporating a wider range of baselines. Specifically, we compare various optimizer choices, including Adam and AdaBelief [Zhuang et al., 2020], alongside MLP architectures such as the standard residuapl MLP [He et al., 2016] and DenseNet [Huang et al., 2017]. These architectures have been previously explored in the context of scaling networks in online deep RL [Lee et al., 2025, Ota et al., 2024], providing a relevant basis for comparison.
We also evaluated state-of-the-art optimizers that have demonstrated success in training large-scale models such as transformers in supervised learning. Specifically, we tested Shampoo [Gupta et al., 2018], a second-order optimizer that maintains and preconditions gradients using full-matrix statistics per layer, and Apollo [Ma, 2020], an adaptive optimizer that leverages curvature information without explicitly computing or storing second-order matrices.
Despite extensive hyperparameter tuning for both methods, we were unable to achieve strong performance in the online deep RL setting. This suggests that further investigation is needed to understand the key properties required for these optimizers to be effective in this regime. Asadi et al. [2023], Ceron and Castro [2021] demonstrate that optimizer behavior plays a critical role in the training dynamics of online deep RL methods, with Asadi et al. [2023] showing that stale optimizer states can hinder learning, and Ceron and Castro [2021] revealing that optimizer sensitivity interacts with the choice of loss function, particularly when comparing Huber and MSE losses.
We present the results for PPO and PQN across all tested optimizers in Fig. 14.
Figure 14: Median human normalized scores on Atari-10 for PPO (top row) and PQN (bottom row), comparing a range of optimizers including RAdam, AdaBelief, Shampoo, Apollo, and Kron (shown in the main curves). While adaptive optimizers like AdaBelief show some robustness, only Kron consistently enables stable and performant training as models scale. Each curve represents the mean performance across three random seeds per algorithm, with shaded areas indicating 95% bootstrap confidence intervals.
C.4 Results with the Multi-Skip Architecture.
We present the full learning curves comparing the proposed multi-skip architecture to the baseline fully connected architecture across all depths and widths studied in the paper. We follow the experimental protocol of Obando Ceron et al. [2023], Ceron et al. [2024c], Agarwal et al. [2021], running each experiment with three random seeds.
Figure 15: Median human-normalized scores with PQN on the Atari-10 benchmark, comparing the baseline agent and the proposed multi-skip architecture across varying depths and widths. The multi-skip architecture not only improves performance at shallow depths, but also enables PQN to remain trainable across all scales considered, whereas the baseline MLP rapidly collapses as depth and width increase. Each curve represents the mean performance across three random seeds per algorithm, with shaded areas indicating 95% bootstrap confidence intervals.
C.5 Ablation on the number of skip connections.
To isolate the effect of skip length on performance, we fix the main network of our proposed MultiSkip architecture (in large size, which includes 5 residual blocks) and vary how many of these blocks receive skip connections from the encoder. When Skip = $k$, we apply the encoder features as skip connections to the first $k$ residual blocks immediately following the encoder, while the remaining (5k)(5-k) blocks operate without direct encoder input. The table below reports human-normalized scores on the Atari-10 benchmark Agarwal et al. [2021].
Table 2: Human-normalized scores on the Atari-10 benchmark, varying the number of residual blocks ($k$) that receive skip connections from the encoder. Performance generally improves as more connections are added.
Environment Skip=1 Skip=2 Skip=3 Skip=4 Skip=5
Amidar-v5 0.20 0.17 0.19 0.20 0.36
BattleZone-v5 0.01 0.67 0.62 0.60 0.69
Bowling-v5 0.07 0.04 0.04 0.08 0.23
DoubleDunk-v5 -2.09 -2.00 -1.55 -1.36 -1.32
Frostbite-v5 0.67 0.70 0.79 0.92 0.88
KungFuMaster-v5 0.93 0.95 0.93 1.12 0.98
NameThisGame-v5 0.79 0.65 0.72 0.85 1.24
Phoenix-v5 0.69 0.68 0.68 0.66 0.66
Qbert-v5 0.84 1.01 1.06 0.91 1.09
Riverraid-v5 0.42 0.44 0.65 0.69 0.99
Aggregate (mean) 0.25 0.33 0.41 0.47 0.58
Performance steadily improves as more skip connections are added, peaking when all 5 blocks are connected. This supports our original design decision to broadcast features to all MultiSkip blocks.
C.6 Results with the Kron Optimizer.
We present the full learning curves comparing the Kron optimizer to the baseline RAdam optimizer originally used in PQN [Gallici et al., 2025], across all depths and widths studied in the paper. We follow the experimental protocol of Obando Ceron et al. [2023], Ceron et al. [2024c], Agarwal et al. [2021], running each experiment with three random seeds.
Figure 16: Median human-normalized scores with PQN on the Atari-10 benchmark, comparing the Kron optimizer to the baseline RAdam optimizer across varying depths and widths. Similar to the multi-skip architecture, Kron not only improves performance at shallow depths, but also enables PQN to remain trainable across all scales considered. In contrast, performance with RAdam rapidly collapses as depth and width increase. Each curve represents the mean performance across thee random seeds per algorithm, with shaded areas indicating 95% bootstrap confidence intervals.
C.7 Justification for Selected Interventions.
The choice of Kronecker-factored optimization and the MultiSkip architecture was the result of a systematic exploration of candidate interventions aimed at improving gradient flow, plasticity, and stability in deep RL. We evaluated a wide range of alternative methods from prior work known to mitigate optimization pathologies [Moalla et al., 2024, Juliani and Ash, 2024].
Our evaluated methods, summarized in Tab. 3, included:
  • Second-order and adaptive optimizers: Apollo [Ma, 2020], Shampoo [Gupta et al., 2018], AdaBelief [Zhuang et al., 2020].
  • Regularization: L2 norm penalties [Kumar et al., 2023], weight clipping, and weight decay [Elsayed et al., 2024].
  • Activation functions: GELU [Hendrycks, 2016] and CReLU [Abbas et al., 2023].
  • Learning rate schedules: Cosine annealing and cyclic schedulers.
  • Learning rate scaling: Multiplying and dividing the default learning rate (2.5e-4) by 10 to compensate for increased network scale.
As shown in the table, none of these interventions consistently improved performance compared to our proposed combination. This motivated our decision to focus on the combination of Kronecker-factored optimization and the multi-skip architecture.
Table 3: Comparison of mean human-normalized scores on Atari-10 for alternative interventions. Our proposed method (Ours) is compared against ablations using only a single alternative intervention. The results (3 seeds per experiment) show that no single alternative consistently matches the performance of our combined approach.
Environment Ours Cosine LR Cyclic LR GELU CReLU L2 Norm Weight Clip Weight Decay LR=2.08e-5 LR=3.00e-3
Amidar-v5 0.355897 -0.001663 0.002276 0.015143 0.016514 0.029293 0.012283 0.013013 0.018848 0.078660
BattleZone-v5 0.694566 0.031297 -0.000287 0.015505 -0.004594 -0.004594 0.035604 0.061446 0.019812 0.575407
Bowling-v5 0.231831 0.050145 -0.008721 0.028706 -0.009811 -0.000727 0.030523 0.051235 -0.086483 0.067587
DoubleDunk-v5 -1.318182 -2.318182 -2.318182 -2.363636 -2.318182 -2.363636 -2.454545 -2.454545 -2.454545 -2.090909
Frostbite-v5 0.881087 -0.001921 0.027708 0.008890 -0.010938 0.004755 0.075372 0.021736 0.008034 0.110036
KungFuMaster-v5 0.975251 -0.011055 0.003181 -0.000823 -0.011500 -0.011278 -0.011500 -0.011055 -0.011055 1.362077
NameThisGame-v5 1.242066 -0.030191 -0.132767 -0.261400 -0.198430 -0.068147 -0.197995 -0.148227 -0.134417 0.747251
Phoenix-v5 0.655141 0.072687 -0.087237 -0.090014 -0.090168 -0.068722 -0.090940 -0.093949 -0.102358 0.411968
Qbert-v5 1.094894 0.013814 0.010052 -0.002833 0.008359 0.006102 0.016259 0.000177 0.001305 0.112282
Riverraid-v5 0.993568 -0.000127 0.345100 0.304636 -0.056846 -0.058494 0.293514 -0.044456 0.057765 0.237365
Appendix D Hyper-parameters
Below, we report the hyperparameters used for each algorithm, which largely follow prior work. For consistency and computational reasons, we relied on default hyper-parameters, despite the known sensitivity of deep RL agents to these choices [Ceron et al., 2024a].
Table 4: PQN Hyperparameters
Hyperparameter Value / Description
Learning rate 2.5e-4
Anneal lr False (no learning rate annealing)
Num envs 128 (parallel environments)
Num steps 32 (steps per rollout per environment)
Gamma 0.99 (discount factor)
Num minibatches 32
Update epochs 2 (policy update epochs)
Max grad norm 10.0 (gradient clipping)
Start e 1.0 (initial exploration rate)
End e 0.005 (final exploration rate)
Exploration fraction 0.10 (exploration annealing fraction)
Q lambda 0.65 (Q($\lambda$) parameter)
Use ln True (use layer normalization)
Activation fn relu (activation function)
Table 5: PPO Hyperparameters
Hyperparameter Value / Description
Learning rate 2.5e-4
Num envs 8
Num steps 128 (steps per rollout per environment)
Anneal lr True (learning rate annealing enabled)
Gamma 0.99 (discount factor)
Gae lambda 0.95 (GAE parameter)
Num minibatches 4
Update epochs 4
Norm adv True (normalize advantages)
Clip coef 0.1 (PPO clipping coefficient)
Clip vloss True (clip value loss)
Ent coef 0.01 (entropy regularization coefficient)
Vf coef 0.5 (value function loss coefficient)
Max grad norm 0.5 (gradient clipping threshold)
Use ln False (no layer normalization)
Activation fn relu (activation function)
Shared cnn True (shared CNN between policy and value networks)
Table 6: PPO Hyperparameters for IsaacGym
Hyperparameter Value / Description
Total timesteps 30,000,000
Learning rate 0.0026
Num envs 4096 (parallel environments)
Num steps 16 (steps per rollout)
Anneal lr False (disable learning rate annealing)
Gamma 0.99 (discount factor)
Gae lambda 0.95 (GAE lambda)
Num minibatches 2
Update epochs 4 (update epochs per PPO iteration)
Norm adv True (normalize advantages)
Clip coef 0.2 (policy clipping coefficient)
Clip vloss False (disable value function clipping)
Ent coef 0.0 (entropy coefficient)
Vf coef 2.0 (value function loss coefficient)
Max grad norm 1.0 (max gradient norm)
Use ln False (no layer normalization)
Activation fn relu (activation function)
Table 7: DQN Hyperparameters
Hyperparameter Value / Description
Learning rate 1e-4
Num envs 1
Buffer size 1,000,000 (replay memory size)
Gamma 0.99 (discount factor)
Tau 1.0 (target network update rate)
Target network frequency 1000 (timesteps per target update)
Batch size 32
Start e 1.0 (initial exploration epsilon)
End e 0.01 (final exploration epsilon)
Exploration fraction 0.10 (fraction of total timesteps for decay)
Learning starts 80,000 (timesteps before training starts)
Train frequency 4 (training frequency)
Use ln False (no layer normalization)
Activation fn relu (activation function)
Table 8: Rainbow Hyperparameters
Hyperparameter Value / Description
Learning rate 6.25e-5
Num envs 1
Buffer size 1,000,000 (replay memory size)
Gamma 0.99 (discount factor)
Tau 1.0 (target network update rate)
Target network frequency 8000 (timesteps per target update)
Batch size 32
Start e 1.0 (initial exploration epsilon)
End e 0.01 (final exploration epsilon)
Exploration fraction 0.10 (fraction of total timesteps for decay)
Learning starts 80,000 (timesteps before training starts)
Train frequency 4 (training frequency)
N step 3 (n-step Q-learning horizon)
Prioritized replay alpha 0.5
Prioritized replay beta 0.4
Prioritized replay eps 1e-6
N atoms 51 (number of atoms in distributional RL)
V min -10 (value distribution lower bound)
V max 10 (value distribution upper bound)
Use ln False (no layer normalization)
Activation fn relu (activation function)
Table 9: Image Classification Hyperparameters (CIFAR-10)
Hyperparameter Value
Batch size 256
Epochs 100
Learning rate 0.00025
Table 10: SAC Hyperparameters
Hyperparameter Value / Description
Critic block type SimBa
Critic num blocks {2, 4, 6, 8}
Critic hidden dim {512, 1024, 1536, 2048}
Target critic momentum ($\tau$) 5e-3
Actor block type SimBa
Actor num blocks {1, 2, 3, 4}
Actor hidden dim {128, 256, 384, 512}
Initial temperature ($\alpha_{0}$) 1e-2
Temperature learning rate 1e-4
Target entropy ($\mathcal{H}^{*}$) $|\mathcal{A}|/2$
Batch size 256
Optimizer {AdamW, Kron}
AdamW’s learning rate 1e-4
Kron’s learning rate 5e-5
Optimizer momentum ($\beta_{1}$, $\beta_{2}$) (0.9, 0.999)
Weight decay ($\lambda$) 1e-2
Discount ($\gamma$) Heuristic
Replay ratio 2
Clipped Double Q False
Table 11: DDPG Hyperparameters
Hyperparameter Value / Description
Critic block type SimBa
Critic num blocks {2, 4, 6, 8}
Critic hidden dim {512, 1024, 1536, 2048}
Critic learning rate 1e-4
Target critic momentum ($\tau$) 5e-3
Actor block type SimBa
Actor num blocks {1, 2, 3, 4}
Actor hidden dim {128, 256, 384, 512}
Actor learning rate 1e-4
Exploration noise $\mathcal{N}(0,{0.1}^{2})$
Batch size 256
Optimizer {AdamW, Kron}
AdamW’s learning rate 1e-4
Kron’s learning rate 5e-5
Optimizer momentum ($\beta_{1}$, $\beta_{2}$) (0.9, 0.999)
Weight decay ($\lambda$) 1e-2
Discount ($\gamma$) Heuristic
Replay ratio 2
Clipped Double Q False
Appendix E Compute Details
All experiments were conducted on a single-GPU setup using an NVIDIA RTX 8000, 12 CPU workers, and 50GB of RAM.
Table 12: Training times across model scales for two optimizers K-FAC shows increased cost as depth and width grow.
Depth Width Optimizer Time
RAdam
Small Small Adam 51m
Small Medium Adam 53m
Small Large Adam 57m
Medium Small Adam 1h 4m
Medium Medium Adam 1h 10m
Medium Large Adam 1h 11m
Large Small Adam 1h 18m
Large Medium Adam 1h 18m
Large Large Adam 1h 27m
Kron
Small Small Kron 1h 59m
Small Medium Kron 2h 27m
Small Large Kron 3h 38m
Medium Small Kron 2h 44m
Medium Medium Kron 3h 32m
Medium Large Kron 5h 59m
Large Small Kron 3h 27m
Large Medium Kron 4h 36m
Large Large Kron 7h 42m
E.1 Results on the Full ALE
In this section, we provide the full training curves corresponding to the aggregated results shown in Sec. 4.3, where we evaluate the performance of the PQN and PPO agents on the full set of environments from the ALE after applying our two proposed gradient interventions. The per-environment learning curves are presented in Fig. 17 for PQN and Fig. 18 for PPO. We follow the experimental protocol of Obando Ceron et al. [2023], Ceron et al. [2024c], Agarwal et al. [2021], running each experiment with three random seeds.
Figure 17: Mean human-normalized score on the full ALE suite, comparing the baseline PQN agent (light curves) with the augmented agent using our combined gradient interventions (dark curves).
Figure 18: Mean human-normalized score on the full ALE suite, comparing the baseline PPO agent (light curves) with the augmented agent using our combined gradient interventions (dark curves).
E.2 Simba on DMC
In this section, we present the full results accompanying the experiments combining Simba [Lee et al., 2025] with our proposed gradient interventions, as introduced in Sec. 4.3. For these experiments, we retain Simba’s original architectural choices but replace the AdamW optimizer with Kron.
We compare Simba using both SAC and DDPG as the underlying RL algorithms. While SAC generally outperforms DDPG, we consistently observe that scaling depth and width, either independently or jointly, leads to a degradation in performance with Simba. However, this degradation is mitigated, and in many cases reversed, when using the Kron optimizer, resulting in improved performance as model capacity increases.
The following figures illustrate these findings:
Figure 19: SAC scaling depth
Figure 20: SAC scaling width
Figure 21: SAC scaling both depth and width
Figure 22: DDPG scaling depth
Figure 23: DDPG scaling width
Figure 24: DDPG scaling both depth and width
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims in the abstract and introduction are grounded in empirical evidence gathered under controlled settings, spanning both supervised and reinforcement learning regimes. We identify and characterize gradient pathologies that emerge with increasing model scale. The proposed gradient-based interventions are evaluated in a modular fashion throughout the paper. Furthermore, the scope of our conclusions is clearly delineated, with detailed descriptions of the specific algorithms and environments studied, ensuring alignment between claims and contributions. While our claims are carefully scoped, our extensive experimentation provides preliminary evidence suggesting that the uncovered phenomena may generalize to broader applications in deep learning.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Sec. 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The paper does not include novel theoretical results, but presents well-established claims in Section 2.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The experimental setup is thoroughly documented in each section presenting results. We specify the number of independent runs, the methods used for aggregating results, and provide detailed descriptions of the algorithms, hyperparameters, and tasks. This level of disclosure ensures that the main experimental findings and conclusions can be independently reproduced.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: Although the data and code will not accompany the submission of this work, they will be released upon publication.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide all essential details of the experimental setup in the main paper, including key training configurations. Additional specifics, such as hyperparameter values and environment settings, are included in the Appendix (Appendix D, Appendix A), ensuring that readers have access to all necessary information to understand and interpret the results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Throughout the paper, we report statistical variability using confidence intervals and standard deviations in training curves, error bars in bar plots, and percentiles in boxplots. We also clearly state the aggregation methods used for each plot (e.g., mean, median, inter-quantile mean), ensuring transparency and appropriate interpretation of the statistical significance of our results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: This information is provided in Appendix E.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authors preserved anonymity during the submission of this work.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See Appendix LABEL:app:broader_impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: The paper does not use existing assets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.